# Local Identification in Instrumental Variable Multivariate Quantile Regression Models

### v3

@@ Abstract
In the instrumental variable quantile regression (IVQR) model of [chernozhukov2005iv], a one-dimensional unobserved rank variable monotonically determines a single potential outcome. Even when multiple outcomes are simultaneously of interest, it is common to apply the IVQR model to each of them separately. This practice implicitly assumes that the rank variable of each regression model affects only the corresponding outcome and does not affect the other outcomes. In reality, however, it is often the case that all rank variables together determine the outcomes, which leads to a systematic correlation between the outcomes. To deal with this, we propose a nonlinear IV model that allows for multivariate unobserved heterogeneity, each of which is considered as a rank variable for an observed outcome. We show that the structural function of our model is locally identified under the assumption that the IV and the treatment variable are sufficiently positively correlated.
@@ 1 Introduction
The estimation of quantiles of potential outcomes in the presence of endogeneity has received increasing attention. For this purpose, the instrumental variable quantile regression (IVQR) model introduced by [chernozhukov2005iv] is popular in applications. While the IVQR model assumes that outcome variables are one-dimensional, researchers are often interested in multiple outcomes at the same time. For example, there is much work that studies Engel curves to investigate how the shares of food and leisure expenditures change as the total expenditure changes using the gross earnings of the head of household as an instrument. See for example [blundell2007semi], and [imbens2009identification]. In such cases, researchers typically run IV regression models separately for each dimension of the outcome variable. However, this approach can be problematic for several reasons.
First of all, it ignores how different outcome variables are correlated with each other, and consequently it is silent on the joint distribution of them. In the Engel curve example, the expenditure shares of food and leisure are likely correlated, not only because there is substitutability, i.e., more food, less leisure, but also because their sum must not exceed one.
More importantly, applying the standard quantile model separately makes little sense in many cases because one of the crucial assumptions of [chernozhukov2005iv], the rank similarity, is unrealistic. Specifically, in order to apply [chernozhukov2005iv] to the Engel curve dataset, it must be assumed that individuals choose the expenditure share on food based only on their preference for food. However, this is unlikely to be true, because they should also care about their preference for leisure, which determines how much they can spend on food.
@@
To address these issues, we propose a novel nonlinear model that allows for multidimensional outcome variables that are potentially correlated with each other. We consider a multivariate extension of the potential outcome framework of [chernozhukov2005iv]. For a treatment let be the corresponding -dimensional potential outcome. We assume that conditional on covariates, it is represented as where is a structural function and is a -dimensional rank variable. The rank variable is a normalized random variable and accounts for the heterogeneity of potential outcomes among observationally identical individuals. It can be interpreted as representing unobserved individual characteristics.
Existing work on endogenous quantile models imposes various monotonicity assumptions on the structural function for identification. Generalizing the assumption of [chernozhukov2005iv] that the structural function is the quantile function of the potential outcome, we assume that the derivative of the structural function is symmetric and positive definite. Under this restriction, an increase in the th entry of leads to an increase in the th entry of In addition, an increase in a particular entry of can affect other dimensions of both positively and negatively. This allows the model to capture substitutability/complementarity between different dimensions of the outcome. Running the standard IVQR separately does not have this flexibility.
A vector-valued function with a symmetric and positive definite derivative is characterized by the cyclical monotonicity, which is a notion of monotonicity for vector-valued functions. There are many economically important examples of cyclically monotone functions in contract theory ([rochet1987necessary], [kushnir2021monotone]), revealed preference theory ([ekeland2013housing]), and econometrics ([shi2018estimating], [chernozhukov2021identification]). We show in Section 2.2 that under a reasonable setup, the vector of the optimal expenditure shares of goods is written as such.
In the optimal transport literature, a vector-valued function with a symmetric and positive definite derivative is known to minimize the quadratic transportation cost and called an optimal transport map. Moreover, recent developments in statistical theory have revealed that it has many desirable properties that qualify it as an extension of the quantile function to random vectors (see, e.g., [ekeland2012comonotonic], [chernozhukov2017monge], [hallin2021distribution], [ghosal2022multivariate]). This justifies calling our model an IV multivariate quantile regression model.
@@
Regarding identification, [chernozhukov2005iv] show that in order to identify quantiles of one-dimensional potential outcomes for binary treatments, an IV with binary support satisfying what they call the full rank condition is necessary. For -dimensional potential outcomes, it may seem natural that the support size of an IV needed for the identification is greater than two because there are functions to be identified. Surprisingly, however, our main result, Theorem 3.1, states that an IV with only binary support can achieve the desired identification. At the cost of multidimensionality, the IV is required to be correlated with the treatment status sufficiently positively, which is a stronger condition than the full rank condition, and it becomes more demanding as the dimension of the potential outcomes increases.
@@
Related literature. Our model stands on the IVQR model that is proposed and developed by [chernozhukov2005iv], [chernozhukov2006instrumental], [chernozhukov2007instrumental] and [chernozhukov2013quantile]. Unlike these work, we allow for the multidimensionality in the outcome and unobservable heterogeneity. There are different studies that deal with endogeneity when outcome variables are one-dimensional. [abadie2002instrumental] consider quantile treatment effects based on the local average treatment effect framework developed by [imbens1994identification]. Their approach identifies treatment effects on compliers and is valid only for binary treatments. Extending [chesher2003identification], [imbens2009identification] take the control function approach under the assumption that the selection equation is monotone and the instrumental variable is independent of unobservable disturbances. [gunsilius2023condition] extends the control function approach to the multidimensional outcome setting. [matzkin2008identification] considers nonparametric identification in simultaneous equation models. In the paper, endogenous variables are assumed to be continuous, and the number of them must be equal to the dimension of unobserved heterogeneity, whereas our model allows for discrete treatments and arbitrary randomness in the treatment assignment stage. In the simultaneous equation framework, [blundell2017individual] propose a nonseparable model when a proxy variable of unobservable heterogeneity is available.
Our assumption on the structural function requires it to be a multivariate quantile function that is defined and developed by [ekeland2012comonotonic], [chernozhukov2017monge], [hallin2021distribution], and [ghosal2022multivariate]. Multivariate quantile functions are defined based on the optimal transport theory ([villani2003topics], [villani2009optimal]). [galichon2018optimal] provides a thorough review on economic applications of optimal transportation. Mathematical results from optimal transport theory are used to prove the identification.
@@
Organization: This paper is organized as follows. Section 2 introduces the IV multivariate quantile regression model. Section 3 formally states the local identification result for the model. Section 4 discusses the cost of taking multidimensional potential outcomes into account and outlines the proof of our main theorem. All proofs and mathematical preliminaries are provided in the Appendix.
@@ 2.1 Model
Our model extends [chernozhukov2005iv] so that it accounts for multidimensional outcome vectors. The following is the primitives.
@@
For each it holds that and there exists a function such that is continuously differentiable in the first argument and its derivative is symmetric and positive definite on
For each is independent of conditional on 111This is slightly redundant. It is enough to assume that the conditional independence between and for some but we keep the stronger assumption following the literature.
@@
We call this model an IV multivariate quantile regression model. The use of “multivariate quantile” follows the literature on the extension of classical quantile functions to multivariate random variables. Section 2.4 discusses this point in more detail.
@@
The difference from [chernozhukov2005iv] appears in (A1). It follows from Theorem A.1 that for any reference measure there exists with a symmetric and positive semi-definite derivative such that but it is not necessarily regular as specified in (A1). Assumption (A1) assumes the smoothness of and the strict positive definiteness of its derivative. This assumption holds if both and the distribution of conditional on are sufficiently regular, according to the regularity theory of optimal transport (see, for example, [caffarelli1992regularity]). Hence, for any other sufficiently regular reference measure condition (A1) is satisfied with some and This implies that any causal parameter defined as a functional of the distributions of the potential outcomes, such as average/quantile treatment effects, is independent of the choice of the reference measure. Also, the continuity of imposed in (A1) implies that the range of is compact and so is the support of conditional on
@@
Condition (A2) implies that the IV is independent of potential outcomes conditional on covariates. Condition (A3) is a weak restriction that allows a broad class of assignment rules of treatments. In particular, in (A3) can be correlated with potential outcomes. Condition (A4) is called the rank similarity, which is usually assumed in the literature. This is weaker than the rank invariance assumption that does not change with assigned treatments. For more details of these assumptions, see [chernozhukov2005iv] and [chernozhukov2013quantile].
@@
The unobservable random variable accounts for the heterogeneity of outcomes among the observationally identical individuals. It is interpreted as a rank variable in the sense that the th entry of is monotonically determined by the th entry of conditional on covariates Indeed, by the positive definiteness of the (Assumption (A1)), it holds that where and On top of that, unlike the one-dimensional model, it is also possible that the th entry of affects the th entry of for in a flexible way. More precisely, the partial derivative can be either positive or negative as long as is positive definite, which allows for the correlation between different dimensions of the observed outcome vector.
Technically, one can choose any absolutely continuous probability measure on any compact convex as the reference measure but taking and seems natural in the examples considered in this paper. One could also think of the -dimensional unit ball as and the spherical uniform distribution as as considered in [chernozhukov2017monge] and [hallin2021distribution].
@@
As the LHS of (1) is determined by the joint distribution of the observable variables, the equation gives a conditional moment restriction. That is, any candidate structural function necessarily satisfies (1) in which is replaced by For identification, we consider conditions under which is the only functions that satisfy the equation in the latter section.
@@
For the notation simplicity, let by suppressing covariates and all the following analysis should be understood as conditional on This treatment does not affect identification results because the distribution of observables conditional on is identifiable.
@@
When outcomes are univariate and the reference distribution is the uniform distribution on the unit interval, as in [chernozhukov2005iv], is the -quantile of and satisfies
@@
for all which corresponds to setting and in (1). If the treatment variable is supported on a finite set , this is a simultaneous equation with unknown variables Hence, if the instrument takes on more than or equal to values and the equation system is non-degenerate, it is expected that the solution is unique, at least locally. Indeed, [chernozhukov2005iv] show that this observation is correct. Checking the condition for each quantile level establishes the nonparametric identification of the structural function
@@
This approach works in dimension one because for each the values are determined in the conditional moment restriction (1) by choosing a specific set Unfortunately, it does not apply to higher dimensions. To see this, for consider and fix Then, any set does not make (1) an equation of only, because the image can be fully nonlinear. Specifically, the image set depends on the values of not only at but also at all the boundary points of Since the set is an infinite-dimensional object, considering (1) for fixed separately does not provide any identification power with a finitely supported IV. To overcome this difficulty, we instead think of (1) as a measure-valued equation in order to exploit the relationship between and for different sets which is not necessary for the identification of univariate structural functions.
2.2 Example: the Engel curve
We illustrate an example the multivariate IV quantile model naturally fits. We consider the Engel curve and investigate how the shares of expenditures on goods change as the total expenditure varies using the gross earnings of the head of household as an instrument.
Suppose we are interested in two goods with price We combine all the other goods into one and call it that has price These prices are assumed to be fixed. A consumer earns He decides his consumption schedule His utility from consuming and paying is where is a concave function and represents his “taste” about the goods. The first two terms are obtained by decomposing a concave utility function into the linear part and the other This formulation assumes that consumers are heterogeneous in the coefficients of the linear component of the utility function. He can choose to save instead of consuming. By saving he receives a (discounted) value Thus, he solves the following maximization problem:
where This is rewritten as
where the new variable represents the total expenditure. This formulation reflects the fact that people typically determine how much to save first and then allocate the budget to the goods. Let be the maximizer of the outer maximization problem and be that of the inner one at For fixed expenditure level we denote the optimal expenditure share of good by
As the following proposition claims, the optimal share depends only on a transformed preference measure where for The vector can be thought of as a relative propensity to consume compared to other goods in the consumption basket.
Proposition 2.1.
Assume and Then, given total expenditure level the optimal expenditure share vector is a function of and written as where is a convex function In particular, if is differentiable, its derivative is symmetric and positive semi-definite.
Suppose that consumers are randomly given the gross earnings and the preference vector which are assumed to be independent. They maximize (2) and find the optimal total expenditure and shares where is associated with If an econometrician observes but not and then this setup is consistent with the IV multivariate quantile regression model.
2.3 Comparison with [chernozhukov2005iv]
Even when several outcome variables are present, the one-dimensional quantile model of [chernozhukov2005iv] could be applied to each dimension. However, this approach has some problems. First of all, it does not account for correlation between different outcome variables. In the Engel curve example in the previous subsection, the expenditure shares of food and leisure are correlated, because their sum must be no more than one.

### v4

@@ Abstract
In the instrumental variable quantile regression (IVQR) model of [12], a one-dimensional unobserved rank variable monotonically determines a single potential outcome. In practice, when researchers are interested in multiple outcomes, it is common to estimate separate IVQR models for each of them. This approach implicitly assumes that the rank variable in each regression affects only its associated outcome, without influencing others. In reality, however, outcomes are often jointly determined by multiple latent factors, inducing structural correlations across equations.
To address this limitation, we propose a nonlinear instrumental variable model that accommodates multivariate unobserved heterogeneity, where each component of the latent vector acts as a rank variable corresponding to an observed outcome. When both the treatment and the instrument are discrete, we show that the structural function in our model is locally identified under a sufficiently strong positive correlation between the treatment and the instrument.
@@ 1 Introduction
The instrumental variable quantile regression (IVQR) model introduced by [12] has become a widely used tool for estimating quantiles of potential outcomes in the presence of endogeneity. See, for example, [11] and [2]. While the IVQR framework assumes a one-dimensional outcome variable, researchers are often interested in settings with multiple outcomes. For instance, [11] examine the effect of 401(k) participation on several wealth measures, including total wealth and financial assets. In such cases, it is common to estimate separate IVQR models for each outcome dimension. However, this practice raises several conceptual and empirical concerns.
First, running separate models ignores the correlation structure among outcome variables and is therefore silent about their joint distribution. In the 401(k) example, this approach cannot capture the share of total wealth accounted for by financial assets.
More importantly, estimating separate quantile models makes rank similarity, a crucial assumption of [12], unrealistic. To apply the IVQR model in the 401(k) example, one must assume that individuals’ holdings of financial assets depend solely on their preferences for those assets. In reality, individuals allocate wealth across multiple assets simultaneously, and their preferences over financial assets are inherently linked to preferences over other asset categories.
@@
To address these limitations, we propose a new nonlinear model that accommodates multidimensional outcomes that may be correlated with each other. Specifically, we consider a multivariate extension of the potential outcome framework in [12]. For a treatment , let denote the corresponding -dimensional potential outcome. Conditional on covariates, we assume it can be represented as , where is a structural function and is a -dimensional rank variable. The rank variable is a random vector that captures unobserved heterogeneity among observationally identical individuals, and can be interpreted as reflecting latent individual characteristics.
Existing studies of endogenous quantile models impose various monotonicity restrictions on the structural function for identification. Extending the assumption in [12] that the structural function is the quantile function of the potential outcome, we require its derivative to be symmetric and positive definite. This restriction implies that an increase in the th component of increases the th component of , while also allowing cross-dimensional effects: changes in one component of can affect other components of positively or negatively. Thus, our framework can capture substitutability and complementarity between outcome dimensions—features that the standard IVQR approach cannot accommodate. Section 2.2 presents two examples of structural functions satisfying this symmetry and positive definiteness condition.
Vector-valued functions with symmetric and positive definite derivatives play a central role in optimal transport theory, where they characterize maps minimizing quadratic transportation costs. Recent developments in statistics have shown that such maps possess many desirable properties that justify viewing them as multivariate analogues of quantile functions (see, e.g., [21], [9], [28], [26]). This perspective motivates us to refer to our framework as an instrumental variable multivariate quantile regression model.
@@
For discrete treatments and instruments, our identification result generalizes the key insight of [12]. They show that to identify quantiles of one-dimensional potential outcomes under binary treatment, an instrument with binary or richer support satisfying the full-rank condition is required. For higher-dimensional potential outcomes, one might expect that instruments with more than two support points are necessary. Surprisingly, our main result, Theorem 3.1, shows that an instrument with only binary support suffices for identification. However, this comes at the cost of a stronger relevance condition: the instrument must be sufficiently positively correlated with the treatment, and the strength required increases with the dimension of the outcome vector.
Related literature. Our model builds on the IVQR framework proposed and developed by [12], [13], [15], and [14]. Unlike these studies, we allow for multidimensional outcomes and multidimensional unobserved heterogeneity.
@@
Several related papers consider endogeneity in models with a single outcome variable. [1] analyze quantile treatment effects using the local average treatment effect framework of [29]. Their approach identifies treatment effects for compliers but is limited to binary treatments. Extending [16], [30] adopt a control function approach under the assumption that the selection equation is monotone and the instrumental variable is independent of the unobserved disturbance. Nonparametric identification in triangular models with discrete instruments is studied by [37], and further extended to multivariate settings by [27]. [31] investigates nonparametric identification in simultaneous equation models, where endogenous variables are continuous and their dimension must coincide with that of unobserved heterogeneity. In contrast, our framework accommodates discrete treatments and allows for arbitrary randomness in the treatment assignment process. Within the same simultaneous equation framework, [3] propose a nonseparable model that exploits proxy variables for unobserved heterogeneity.
Our key restriction on the structural function requires it to be a multivariate quantile function, as developed by [21], [9], [28], and [26]. Multivariate quantile functions are grounded in optimal transport theory ([39], [40]), which provides a natural framework for extending quantiles to random vectors. [25] offers a comprehensive overview of economic applications of optimal transport. In this paper, we draw on results from this literature to establish the identification of our model.
@@
Organization: This paper is organized as follows. Section 2 introduces the IV multivariate quantile regression model. Section 3 formally states the local identification result for the model and discusses the relationship to existing results. Section 4 concludes. All proofs and mathematical preliminaries are provided in the Appendix.
@@ 2.1 Model
Our model extends [12] so that it accounts for multidimensional outcome vectors. The following is the primitives.
@@
For each it holds that and there exists a function such that is continuously differentiable in the first variable, and its derivative is symmetric and positive definite on
For each is independent of conditional on
@@
We refer to this framework as an IV multivariate quantile regression model. The term “multivariate quantile” follows the literature that extends the classical notion of quantile functions to multivariate random variables. Further details are provided in Appendix A. See also, for example, [6], [9], [28], and [26].
@@
The difference from [12] appears in (A1). It follows from Theorem A.1 that for any reference measure there exists with a symmetric and positive semi-definite derivative such that but it is not necessarily regular as specified in (A1). Assumption (A1) assumes the smoothness of and the strict positive definiteness of its derivative. This assumption holds if both and the distribution of conditional on are sufficiently regular, according to the regularity theory of optimal transport (see, for example, [5]). Hence, for any other sufficiently regular reference measure condition (A1) is satisfied with some and This implies that any causal parameter defined as a functional of the distributions of the potential outcomes, such as average/quantile treatment effects, is independent of the choice of the reference measure. Also, the continuity of imposed in (A1) implies that the range of is compact and so is the support of conditional on
@@
Condition (A2) implies that the IV is independent of potential outcomes conditional on covariates. Condition (A3) is a weak restriction that allows a broad class of assignment rules of treatments. In particular, in (A3) can be correlated with potential outcomes.
Condition (A4) is called the rank similarity ([12]). The simplest form of rank similarity is rank invariance that requires for all In the wealth accumulation example discussed in Section 1 and 2.2, the rank invariance assumes that an individual’s preference does not change regardless of their participation to 401(k). On the other hand, the rank similarity assumes that the preference may change depending on the participation status, but the individual cannot predict the change before deciding whether to participate 401(k) or not. For more details of these assumptions, see [12] and [14].
One limitation of the IV multivariate quantile regression model is that Condition (A4) depends on the choice of the reference measure . In particular, the fact that Condition (A4) holds for a given does not guarantee that it will hold for another measure . This dependence implies that the distribution of unobserved heterogeneity must be specified a priori. Although strong, such an assumption is common in related literature. For example, [10] impose a similar condition for identifying hedonic equilibrium models, and also, discrete-choice models typically rely on a fixed distribution for unobserved heterogeneity.
@@
The unobserved random vector captures heterogeneity in outcomes among observationally identical individuals. It is interpreted as a rank variable in the sense that, conditional on covariates the th component of is monotonically determined by the corresponding component of In particular, under the positive definiteness of imposed in Assumption (A1), we have , where and . Moreover, unlike in the one-dimensional setting, the th component of may also influence the th component of for . Specifically, the cross-partial derivative may be either positive or negative, as long as remains positive definite. This feature enables the model to capture flexible correlations and potential substitutability or complementarity across different dimensions of the outcome vector.
Technically, any absolutely continuous probability measure on a compact convex set can be chosen as the reference measure . In the absence of a strong reason to prefer a particular specification, it is natural to take and , as in [6]. Alternatively, one may consider the -dimensional unit ball as and the spherical uniform distribution as , following [9] and [28].
@@
As the LHS of (1) is determined by the joint distribution of the observable variables, the equation gives a conditional moment restriction. In particular, any candidate structural function must satisfy (1) with replaced by . For identification, we explore conditions under which is the only function that satisfies the equation in the subsequent section.
@@
For notational simplicity, let , suppressing the dependence on covariates . All subsequent analysis should be understood as conditional on . This simplification does not affect the identification results, since the conditional distribution of the observables given is identifiable.
@@
When outcomes are univariate and the reference distribution is the uniform distribution on the unit interval, as in [12], is the -quantile of and satisfies
@@
for all which corresponds to setting and in (1). If the treatment variable is supported on a finite set , this is a simultaneous equation with unknown variables Hence, if the instrument takes on more than or equal to values and the equation system is non-degenerate, it is expected that the solution is unique, at least locally. Indeed, [12] show that this observation is correct. Checking the condition for each quantile level establishes the nonparametric identification of the structural function
@@
This approach works in one dimension because, for each , the values are determined by the conditional moment restriction (1) when we take the set . In higher dimensions, however, this argument no longer applies. To see why, consider the case with and fix . For any set , equation (1) does not isolate alone, because the image can be fully nonlinear. In particular, the image depends not only on the value of at but also on its behavior along the boundary . Since the set is an infinite-dimensional object, evaluating (1) for a fixed does not yield identification when the instrument has finite support. To address this issue, we instead interpret (1) as a measure-valued equation, allowing us to exploit the relationships between and for different sets and . Such relationships are unnecessary in the univariate case but are crucial for identifying multivariate structural functions.
2.2 Examples
Example 1. We first consider an extension of the saving model proposed by [11]. The paper investigates the effects of 401(k) participation on wealth accumulation. As measures of wealth, they consider total wealth, net financial assets, and net non-401(k) financial assets. Since [11] apply the IVQR model to these variables separately, their analysis does not capture the correlation among them. For example, while total wealth is defined as net financial assets plus other assets (e.g., housing equity and the value of business, property, and motor vehicles), their model has no implication for how people distribute total wealth into these sub categories. We shall see that our multivariate model can take both variables into consideration simultaneously and allows us to discuss their joint distribution.
Let represent the preference for net financial assets and other assets under participation status Let be net financial assets and other assets, respectively. An individual with type whose portfolio is receives a quasi-linear utility where is a deterministic concave function.111Notice that covariates are suppressed for simplicity. Assuming an interior solution, the individual chooses the optimal portfolio which follows by the first order condition. It can be seen that the function has a symmetric and positive definite derivative by the concavity of
The individual determines whether to participate 401(k) based on the eligibility of 401(k) and unobservable heterogeneity This is represented by The eligibility variable is considered to work as a valid IV. The joint distribution of fits into the multivariate IVQR model under and the rank invariance/similarity.
With identified, we can obtain the joint distribution of net financial assets and other assets. This allows us to discuss the effect of 401(k) participation on the portfolio selection, which is impossible in the standard IVQR framework.
Example 2. We now show how a simple discrete choice problem fits naturally into the IV multivariate quantile framework. This example is inspired by [35] and [24]. Consider a retailer offering differentiated goods in two sets of stores, indexed by The products’ qualities are captured by a latent quality vector where is assumed known (e.g., from in-store audits of freshness and display quality). Each consumer in district has an unobservable preference vector The preference distributions are unknown and may differ because district is exposed to a marketing campaign. By choosing product a consumer receives utility
The analyst never observes or . Instead, one observes the aggregate market share vector if the store is in district where
Economically, represents how a marginal improvement in the quality redistributes sales across all goods under the taste regime .
The Williams-Daly-Zachary theorem implies that the surplus function defined as
is related to the market share vector via the equation (see, e.g., [33]). It is not hard to see that is a convex function. Also, it is differentiable under mild regularity on Hence, has a symmetric and positive definite derivative.
Let denote the district, or equivalently the treatment status, ( or ) and suppose there exists a valid instrument (e.g. randomized ad‐exposure eligibility). Then we observe i.i.d. draws of with and This setup satisfies the rank invariance and exogeneity conditions of the IV multivariate quantile model.
Applying a separate IVQR mode of [12] to each product discards the rich cross‐product relationships. For instance, one-dimensional approaches cannot impose the natural constraint nor capture correlations in arising from similar products facing related quality shocks. The multivariate model preserves these economically meaningful cross‐product patterns while identifying the two demand-quality mappings and
2.3 Comparison with [12]